### v2

# Chirped Gaussians Have Maximal Frame Sets

Abstract
Let $\varphi(x)=e^{-\pi x^{2}}$ be the Gaussian and $h_{\lambda}(x)=e^{-\pi i\lambda x^{2}}$ be a chirp where $\lambda\in\mathbb{R}\setminus\{0\}$ is a parameter. For $\gamma>0$, let $\varphi_{\gamma}(x)=e^{-\pi\gamma^{2}x^{2}}$ be the dilated Gaussian, we prove that for any such $\lambda,\gamma$, the chirped Gaussian $h_{\lambda}\cdot\varphi_{\gamma}$ always has maximal frame set, i.e., its frame set consists of precisely all positive pairs $(\alpha,\beta)$ with $\alpha\beta<1$. The proof is by using fractional Fourier transform to establish maximality on certain product-convoluted (with chirps) Gaussians first, then reduce general single chirped cases to it. It follows that $\mathcal{G}(\varphi_{\gamma},Q\mathbb{Z}^{2})$ is a frame for $Q\in\mathbb{R}^{2\times 2}$ if and only if $0<|\det Q|<1$. In addition, with the theta function we also show that the Zak transform $Z(h_{\lambda}\cdot\varphi_{\gamma})$ always has a unique simple zero at the center of the unit square.
Keywords: Gabor frames; Gaussians; chirps; Totally positive functions; fractional Fourier transforms; Zak transforms
2020 MSC: 42C15; 42C40
1 Introduction
Let
$T_{t}:f(x)\mapsto f(x-t),\quad M_{\xi}:f(x)\mapsto e^{2\pi i\xi x}f(x),$
be respectively the translation (i.e., time shift) operator and the modulation (i.e., frequency shift) operator on $L^{2}(\mathbb{R})$. Given $g\in L^{2}(\mathbb{R})$ and $\Lambda\subset\mathbb{R}^{2}$, a Gabor system on $L^{2}(\mathbb{R})$ with window $g$ and support $\Lambda$ is defined and denoted by
$\mathcal{G}(g,\Lambda)=\{M_{a}T_{b}g:(b,a)\in\Lambda\}.$
A typical choice of $\Lambda$ would be the lattice $\alpha\mathbb{Z}\times\beta\mathbb{Z}$ with $\alpha,\beta>0$. And the corresponding Gabor system is
$\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})=\{M_{\beta j}T_{\alpha k}g:j,k\in\mathbb{Z}\}.$
Gabor systems are fundamental building blocks in signal processing and time-frequency analysis since they offer a natural way to discretize a function on the time-frequency plane. A trivial example of a Gabor system is the orthonormal basis of $L^{2}(\mathbb{R})$ formed by taking the window to be the characteristic function on the unit interval $\chi_{[0,1)}$ and $\alpha=\beta=1$.
A set of functions $\{h_{k}\}_{k\in\mathbb{Z}}$ is called a frame in a Hilbert space $H$ if there exist constants $A,B>0$ so that
$A\|f\|_{H}^{2}\leq\sum_{k\in\mathbb{Z}}|\langle f,h_{k}\rangle_{H}|^{2}\leq B\|f\|_{H}^{2}$
holds for any $f\in H$. The constants $A,B$ are then called respectively the lower and the upper frame constants of this frame. If $\{h_{k}\}_{k\in\mathbb{Z}}$ is a frame, then both the analysis operator $f\mapsto\{\langle f,h_{k}\rangle_{H}\}_{k\in\mathbb{Z}}$ and the synthesis operator $\{\langle f,h_{k}\rangle_{H}\}_{k\in\mathbb{Z}}\mapsto f$ are bounded from above and below. Hence it allows a stable decomposition and reconstruction of $f$ into and from the coefficients $\{\langle f,h_{k}\rangle_{H}\}_{k\in\mathbb{Z}}$.
A Gabor frame on $L^{2}(\mathbb{R})$ is a Gabor system that forms a frame on $L^{2}(\mathbb{R})$. The set of parameters $(\alpha,\beta)$ that makes $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ a frame is called the frame set of $g$ and denoted by $F(g)$. The problem of determining $F(g)$ for an arbitrary function $g$ is a long standing difficult open problem in this field. Even for simple window functions such as $\chi_{[0,1)}$, the corresponding frame set could be fairly complicated (see [14]). The classical density theory points out that if $\alpha\beta>1$, then $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is incomplete and thus can not possibly be a frame (see e.g., [10, Theorem 3a] or [1]).
The critical case $\alpha\beta=1$ is in general not so interesting since such windows are already completely characterized by the Zak transform. More importantly it has been excluded from frame sets of functions with good time-frequency concentration. This is due to the fact that in such case $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is a frame if and only if the Zak transform of $g$ is essentially bounded away from both 0 and $\infty$ (see e.g., [6, Corollary 8.3.2]), and it is well known that if the transformed function is continuous, then it always admits zeros (see e.g., [6, Lemma 8.4.2]) and thus can not be lowered bounded. A class of functions that will guarantee continuity after the Zak transform is the set of continuous Wiener functions. Wiener functions on $\mathbb{R}$ form a Banach space equipped with the norm (see e.g., [6, Chapter 6.1]).
$\|g\|_{W(\mathbb{R})}=\sum_{k\in\mathbb{Z}}\operatorname*{ess\,sup}_{x\in[0,1)}|g(x+k)|<\infty.$
It is thus a mixed norm space that is locally of $L^{\infty}$ bound and globally of $\ell^{1}$ decay. The subspace of continuous functions in it is denoted by $W_{0}(\mathbb{R})$. That the frame set of a $W_{0}(\mathbb{R})$ function (or its Fourier transform) can not contain any pair $(\alpha,\beta)$ with $\alpha\beta=1$ is known as the Amalgam Balian-Low theorem.
We shall say that a function $g$ has maximal frame set (or $F(g)$ is maximal), if $F(g)$ contains all pairs $(\alpha,\beta)$ with $\alpha\beta<1$ and $\alpha,\beta>0$. A primal question is then to identify all $W_{0}(\mathbb{R})$ functions whose frame sets are maximal. It was first conjectured in [4] that Gaussians possess such a property. This is proved in [20, 24, 25]. Later the hyperbolic secant $1/(e^{x}+e^{-x})$ [15], the one sided exponential $e^{-x}\chi_{\mathbb{R}^{+}}(x)$ [12] (actually this is not a $W_{0}(\mathbb{R})$ function since it is not even continuous, its Zak transform has no zero and its frame set contains also the critical case $\alpha\beta=1$, see [14, Section 3.2.1]) and the two sided exponential $e^{-|x|}$ [13] are also found to have this property. Then apparently automorphisms of the $L^{2}$ space such as translations, modulations, dilations, scalings and Fourier transforms of these functions will inherit this property.
It has been noticed that these functions are not only positive but also totally positive. A function $g$ is called totally positive if for all $n\in\mathbb{N}$ and any two sequences of non-decreasing numbers $x_{1}\leq\ldots\leq x_{n}$ and $y_{1}\leq\ldots\leq y_{n}$, the determinant of the matrix $g(x_{i}-y_{j})]_{i,j=1}^{n}$ is non-negative. Alternative $g$ is totally positive if and only if its Fourier transform factors into
$\hat{g}(\xi)=ce^{-\gamma\xi^{2}}e^{2\pi i\nu\xi}\prod_{j=1}^{N}(1+2\pi i\nu_{j}\xi)e^{-2\pi i\nu_{j}\xi},$ (1)
where $c>0$, $\gamma\geq 0$, $\nu,\nu_{j}\in\mathbb{R}$, $N\in\mathbb{N}\cup\{\infty\}$ and $0<\gamma+\sum_{j=1}^{N}\nu_{j}^{2}<\infty$. This is characterized in [21, 22, 23]. In a series of papers [7, 8, 9] it is proved that if $N$ is finite or if $N=\infty$ but $\alpha\beta\in\mathbb{Q}$, then $\alpha\beta<1$ is in the frame set of $g$ (in particular, if $g$ is at the same time in the Feichtinger algebra, then $F(g)$ is open [5], and thereof maximal), which covers and also greatly extends all aforementioned known cases. Mysteriously the Zak transforms of all these functions except the one-sided exponential possess a unique simple zero in the unit square (see [2, 17, 18, 26]).
Is total positivity necessary for maximality of frame sets? If we modulate a totally positive function, or simply multiply it by a negative number, then it is not totally positive any more but its frame set stays invariant. Therefore to discuss necessity it perhaps makes sense to phrase it as “is total positivity necessary for real non-negative Wiener functions to have maximal frame sets?” In fact it is pointed out in [3] that certain linear combinations of Cauchy kernels do admit maximal frame sets, these functions are neither totally positive nor (compositions of) time-frequency shifts or dilations or Fourier transforms of totally positive functions, but they are not Wiener functions either.
The purpose of this paper is to produce a class of complex valued $W_{0}(\mathbb{R})$ functions whose frame sets are maximal and whose Zak transforms also have unique simple zeros at center of the unit square. It is clear from their expressions that these functions are as well neither totally positive nor (compositions of) time-frequency shifts or dilations or Fourier transforms of totally positive functions. Let
$\varphi(x)=e^{-\pi x^{2}},\quad h_{\lambda}(x)=e^{-\pi i\lambda x^{2}}$
be respectively the Gaussian and a chirp where $\lambda\in\mathbb{R}\setminus\{0\}$ is a parameter. For $\gamma>0$, let
$\varphi_{\gamma}(x)=e^{-\pi\gamma^{2}x^{2}}$
be the dilated Gaussian. The main results are the following (Theorem 2, Theorem 3, Theorem 4):
Theorem.
The chirped Gaussian $h_{r}\cdot\varphi_{\sqrt{u}}$ has maximal frame set for any $r\in\mathbb{R}\setminus\{0\}$ and any $u>0$.
Theorem.
Let $\gamma>0$ and $Q\in\mathbb{R}^{2\times 2}$, then $\mathcal{G}(\varphi_{\gamma},Q\mathbb{Z}^{2})$ is a frame if and only if $0<|\det Q|<1$.
Theorem.
$Z(h_{\lambda}\cdot\varphi_{\gamma})$ has a unique simple zero at the center $(1/2,1/2)$ of the unit square $S$ for any $\lambda\in\mathbb{R}\setminus\{0\}$ and any $\gamma>0$.
In the next section we align notations. Since the techniques for maximality of the frame set and for zeros of the Zak transform are quite independent, these results are split into two sections, with the first subsection of each section being an exposition of preliminaries. Theorem 2 and Theorem 3 will be presented in Section 3, and Theorem 4 in the final section. The proof of maximality of $F(h_{r}\cdot\varphi_{\sqrt{u}})$ is by showing in Theorem 1 the maximality of frame sets of product-convoluted (with chirps) Gaussians $h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{1/\gamma})$ for certain $\lambda,\lambda^{\prime},\gamma$ first, then we show that there are $r,u$ so that $h_{r}\cdot\varphi_{\sqrt{u}}=h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{1/\gamma})$ and $r/u$ can take any non-zero real value.
2 Notations and conventions
In this article the Fourier transform of a function $f$ is denoted by $\mathcal{F}(f)$ or $\hat{f}$, and for $f\in L^{1}(\mathbb{R})$ it takes the following integral form:
$\mathcal{F}f(\xi)=\hat{f}(\xi)=\int_{\mathbb{R}}f(x)e^{-2\pi i\xi x}dx.$
The inverse Fourier transform of $f$ is denoted by $\mathcal{F}^{-1}f$.
It is customary and convenient to use the following notation for time-frequency shifts:
$\pi(z)=M_{a}T_{b},$
where $z=(a,b)^{T}\in\mathbb{R}^{2}$ (Hereafter $T$ in the superscript stands for transposition). We also remind the readers about the commutativity relation
$M_{a}T_{b}M_{a^{\prime}}T_{b^{\prime}}=e^{2\pi i(ab^{\prime}-a^{\prime}b)}M_{a^{\prime}}T_{b^{\prime}}M_{a}T_{b}.$ (2)
Let $Q\in\mathbb{C}^{2\times 2}$ and $\mathcal{L}$ a linear operator (could be the identity operator, in which case it is omitted then) on $L^{2}(\mathbb{R})$, to reduce the overuse of the summation symbol, we will adopt the notation $\left\|\langle f,\mathcal{L}\pi(Q\mathbb{Z}^{2})g\rangle\right\|_{\ell^{2}}$ for
$\left\|\langle f,\mathcal{L}\pi(Q\mathbb{Z}^{2})g\rangle\right\|_{\ell^{2}}=\sqrt{\sum_{z\in\mathbb{Z}^{2}}\left|\langle f,\mathcal{L}\pi(Qz)g\rangle\right|^{2}},$
where $\langle\cdot,\cdot\rangle$ is the inner product on $L^{2}(\mathbb{R})$. The left-hand side is indeed the $\ell^{2}$ norm of the the sequence $\{\langle f,\mathcal{L}\pi(Qz)g\rangle\}_{z\in\mathbb{Z}^{2}}$ generated by fixed $f,g,Q,\mathcal{L}$. Taking $Q=\operatorname{diag}(\beta,\alpha)$ and $\mathcal{L}$ the identity operator we get the sequence that corresponds to the usual lattice Gabor system $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$.
We will also use the $\asymp$ symbol for equivalence type inequalities, i.e., we may write
$\left\|\langle f,\pi(Q\mathbb{Z}^{2})g\rangle\right\|_{\ell^{2}}\asymp\left\|f\right\|_{L^{2}(\mathbb{R})},$
instead of $\sqrt{A}\left\|f\right\|_{L^{2}(\mathbb{R})}\leq\left\|\langle f,\pi(Q\mathbb{Z}^{2})g\rangle\right\|_{\ell^{2}}\leq\sqrt{B}\left\|f\right\|_{L^{2}(\mathbb{R})}$.
The dilation of a function $g$ by a parameter $\gamma>0$ will be denote by
$g_{\gamma}(x)=g(\gamma x).$
It is easy to see that
$\left\|g\right\|_{L^{2}(\mathbb{R})}^{2}=\int_{\mathbb{R}}|g(x)|^{2}dx=\gamma\int_{\mathbb{R}}|g(\gamma y)|^{2}dy=\gamma\left\|g_{\gamma}\right\|_{L^{2}(\mathbb{R})}^{2}.$ (3)
Similarly if $D_{\gamma}=\operatorname{diag}(1/\gamma,\gamma)$, then for any $z=(a,b)^{T}\in\mathbb{R}^{2}$
$|\langle f,\pi(D_{\gamma}z)g\rangle|=|\langle f,M_{a/\gamma}T_{b\gamma}z)g\rangle|=\gamma|\langle f_{\gamma},M_{a}T_{b}g_{\gamma}\rangle|=\gamma|\langle f_{\gamma},\pi(z)g_{\gamma}\rangle|.$ (4)
There might be confusions concerning this notation of dilation since $h_{\lambda}(x)=e^{-\pi i\lambda x^{2}}$ is not a dilation of $h_{1}(x)$ by $\lambda$, but we have to write $h_{\lambda}$ this way because we need $\lambda<0$ in certain cases, which can not be achieved by dilations. In compensation for that we will never dilate a chirp in this paper, dilations will only be applied to $L^{2}(\mathbb{R})$ functions.
3 Frame sets of chirped Gaussians
3.1 The fractional Fourier transform and chirps
Let $2\pi\mathbb{T}$ be the Torus obtained by taking $\mathbb{R}$ modulo $2\pi$, and identified as $0,2\pi)$. The fraction Fourier transform $\{\mathcal{F}^{(\theta)}\}_{\theta\in 2\pi\mathbb{T}}$ is a family of operators on $L^{2}(\mathbb{R})$ with $\mathcal{F}^{(\pi/2)}$ being the usual Fourier transform, and satisfying
$\mathcal{F}^{(\theta)}\mathcal{F}^{(\theta^{\prime})}=\mathcal{F}^{(\theta+\theta^{\prime})}.$
It is thus a representation of the additive Abelian group $2\pi\mathbb{T}$ with $\mathcal{F}^{(0)}$ being the identity. For $\theta\notin\pi\mathbb{Z}$ and $f\in L^{1}(\mathbb{R})$ it has the following explicit integral formula:
$\mathcal{F}^{(\theta)}f(\xi)=\sqrt{1-i\cot\theta}\int_{\mathbb{R}}f(x)e^{i\pi\left((x^{2}+\xi^{2})\cot\theta-2x\xi\csc\theta\right)}dx.$
The angle $\theta$ has the meaning that the Wigner distribution of $\mathcal{F}^{(\theta)}f$ differs from that of $f$ by a counterclockwise axis rotation of $\theta$ on the time-frequency plane. This also gives an intuitive explanation to property (ii) below, since the Wigner distribution of $\varphi$ is the two dimensional Gaussian, which is rotational invariant.
Proposition 1.
  1. (i)
    $\mathcal{F}^{(\theta)}$ is unitary on $L^{2}(\mathbb{R})$.
  2. (ii)
    $\mathcal{F}^{(\theta)}\varphi=\varphi$ holds for all $\theta\in 2\pi\mathbb{T}$.
  3. (iii)
    Let $z=(a,b)^{T}\in\mathbb{R}^{2}$ and
    $R_{\theta}=\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix}$, (5)
    then we have
    $\pi(z)\mathcal{F}^{(\theta)}=e^{-2\pi i\left(\frac{(b^{2}-a^{2})\sin(2\theta)}{4}-ab\sin^{2}\theta\right)}\mathcal{F}^{(\theta)}\pi(R_{\theta}z).$
These properties are well known, see perhaps [19, Section 4] and references therein for a concise exposition of properties of $\mathcal{F}^{(\theta)}$ tailored to time-frequency analysis.
For any $f\in L^{2}(\mathbb{R})$, we define $h_{\lambda^{\prime}}*f$ to be
[…]
Lemma 1.
Let
$U_{\lambda}=\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix},\quad L_{\lambda^{\prime}}=\begin{pmatrix}1&0\\ \frac{1}{\lambda^{\prime}}&1\end{pmatrix}$.
Then the following properties hold for any $f\in L^{2}(\mathbb{R})$ and any $z=(a,b)^{T}\in\mathbb{R}^{2}$:
  1. (i)
    $\left\|h_{\lambda}\cdot(h_{\lambda^{\prime}}*f)\right\|_{L^{2}(\mathbb{R})}=\left\|h_{\lambda^{\prime}}*(h_{\lambda}\cdot f)\right\|_{L^{2}(\mathbb{R})}=|\lambda^{\prime}|^{-\frac{1}{2}}\left\|f\right\|_{L^{2}(\mathbb{R})}.$
  2. (ii)
    $\pi(z)(h_{\lambda}\cdot f)=e^{-\pi i\lambda b^{2}}h_{\lambda}\cdot\left(\pi(U_{\lambda}z)f\right).$
  3. (iii)
    $\pi(z)(h_{\lambda^{\prime}}*f)=e^{-\frac{\pi ia^{2}}{\lambda^{\prime}}}h_{\lambda^{\prime}}*\left(\pi(L_{\lambda^{\prime}}z)f\right).$
Proof.
Recall that if $\lambda^{\prime}>0$, then
$\hat{h}_{\lambda^{\prime}}=\frac{1}{\sqrt{\lambda^{\prime}}}e^{-\frac{\pi i}{4}}h_{-\frac{1}{\lambda^{\prime}}},$
where the term $e^{-\pi i/4}$ comes from taking the reciprocal of a square root of $i$, while if $\lambda^{\prime}<0$, then we have to take it on i-i and the formula becomes (alternatively this can also be obtained by applying $\mathcal{F}$ again, see also the table in [16, A-7])
$\hat{h}_{\lambda^{\prime}}=\frac{1}{\sqrt{|\lambda^{\prime}|}}e^{\frac{\pi i}{4}}h_{-\frac{1}{\lambda^{\prime}}}.$
Since the result of our derivations does not depend on whether it is $e^{\pi i/4}$ or $e^{-\pi i/4}$ in the constant $|\lambda^{\prime}|^{-1/2}e^{\pm\pi i/4}$, we simply denote this constant by $c_{\lambda^{\prime}}$ so that
$\hat{h}_{\lambda^{\prime}}=c_{\lambda^{\prime}}h_{-\frac{1}{\lambda^{\prime}}}.$ (6)
Consequently
$\left\|h_{\lambda}\cdot(h_{\lambda^{\prime}}*f)\right\|_{L^{2}(\mathbb{R})}=\left\|h_{\lambda^{\prime}}*f\right\|_{L^{2}(\mathbb{R})}=|c_{\lambda^{\prime}}|\cdot\left\|h_{-\frac{1}{\lambda^{\prime}}}\cdot f\right\|_{L^{2}(\mathbb{R})}=|\lambda^{\prime}|^{-\frac{1}{2}}\left\|f\right\|_{L^{2}(\mathbb{R})}.$
$\left\|h_{\lambda^{\prime}}*(h_{\lambda}\cdot f)\right\|_{L^{2}(\mathbb{R})}$ is obtained in the same way, together these establish (i). Next, straightforward computation shows
$\pi(z)(h_{\lambda}f)$ $=e^{2\pi iax}\cdot e^{-\pi i\lambda(x-b)^{2}}\cdot f(x-b)$
$=e^{2\pi iax}\cdot e^{-\pi i\lambda x^{2}}\cdot e^{2\pi i\lambda bx}\cdot e^{-\pi i\lambda b^{2}}\cdot f(x-b)$
$=e^{-\pi i\lambda b^{2}}\cdot h_{\lambda}\cdot(M_{a+\lambda b}T_{b}f)$
$=e^{-\pi i\lambda b^{2}}h_{\lambda}\cdot\left(\pi(U_{\lambda}z)f\right),$
which is (ii). Finally, with similar computations we get
$\pi(z)(h_{\lambda^{\prime}}*f)$ $=\mathcal{F}^{-1}\mathcal{F}\left(\pi(z)(h_{\lambda^{\prime}}*f)\right)$
[…]
$=e^{2\pi iab}\cdot e^{\frac{\pi ia^{2}}{\lambda^{\prime}}}\cdot h_{\lambda^{\prime}}*(T_{b+\frac{a}{\lambda^{\prime}}}M_{a}f)$ by (6)
$=e^{-\frac{\pi ia^{2}}{\lambda^{\prime}}}h_{\lambda^{\prime}}*(M_{a}T_{\frac{a}{\lambda^{\prime}}+b}f)$ by (2)
$=e^{-\frac{\pi ia^{2}}{\lambda^{\prime}}}h_{\lambda^{\prime}}*\left(\pi(L_{\lambda^{\prime}}z)f\right).$
This establishes (iii). ∎
3.2 Maximality: a special case
Lemma 2.
Let $R_{\theta}$ be as defined in Proposition 1 and
$D_{\gamma}=\operatorname{diag}(\frac{1}{\gamma},\gamma),\quad D_{\alpha,\beta}=\operatorname{diag}(\beta,\alpha).$
Set
$Q_{\gamma,\theta,\alpha,\beta}=D_{\gamma}R_{\theta}D_{\alpha,\beta}=\begin{pmatrix}\frac{1}{\gamma}&0\\ 0&\gamma\end{pmatrix}\begin{pmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}\beta&0\\ 0&\alpha\end{pmatrix}$.
Then $\mathcal{G}(\varphi_{1/\gamma},Q_{\gamma,\theta,\alpha,\beta}\mathbb{Z}^{2})$ is a Gabor frame for any $\gamma>0$, any $\theta\in 2\pi\mathbb{T}$ and any pair $(\alpha,\beta)$ with $\alpha,\beta>0,\alpha\beta<1$.
Proof.
For any $f,g\in L^{2}(\mathbb{R})$, simply letting $(a,b)^{T}$ in (4) run over elements of $R_{\theta}D_{\alpha,\beta}\mathbb{Z}^{2}$ and then summing up we get
$\left\|\langle f,\pi(Q_{\gamma,\theta,\alpha,\beta}\mathbb{Z}^{2})g\rangle\right\|_{\ell^{2}}=\left\|\langle f,\pi(D_{\gamma}R_{\theta}D_{\alpha,\beta}\mathbb{Z}^{2})g\rangle\right\|_{\ell^{2}}\asymp\left\|\langle f_{\gamma},\pi(R_{\theta}D_{\alpha,\beta}\mathbb{Z}^{2})g_{\gamma}\rangle\right\|_{\ell^{2}}.$ (7)
Next, by Proposition 1 (iii) for each $z=(\beta j,\alpha k)^{T}\in D_{\alpha,\beta}\mathbb{Z}^{2}$ we have
$|\langle\mathcal{F}^{(\theta)}f_{\gamma},\mathcal{F}^{(\theta)}\pi(R_{\theta}z)g_{\gamma}\rangle|=|c\langle\mathcal{F}^{(\theta)}f_{\gamma},\pi(z)\mathcal{F}^{(\theta)}g_{\gamma}\rangle|=|\langle\mathcal{F}^{(\theta)}f_{\gamma},\pi(z)\mathcal{F}^{(\theta)}g_{\gamma}\rangle|,$
where (though not import)
$c=e^{-2\pi i\left(\frac{\left((k\alpha)^{2}-(j\beta)^{2}\right)\sin(2\theta)}{4}-jk\alpha\beta\sin^{2}\theta\right)}.$
Summing up we further obtain
$\left\|\langle\mathcal{F}^{(\theta)}f_{\gamma},\mathcal{F}^{(\theta)}\pi(R_{\theta}D_{\alpha,\beta}\mathbb{Z}^{2})g_{\gamma}\rangle\right\|_{\ell^{2}}=\left\|\langle\mathcal{F}^{(\theta)}f_{\gamma},\pi(D_{\alpha,\beta}\mathbb{Z}^{2})\mathcal{F}^{(\theta)}g_{\gamma}\rangle\right\|_{\ell^{2}}.$ (8)
Now we take $g=\varphi_{1/\gamma}$, so that $g_{\gamma}=\varphi$ to get
$\left\|\langle f,\pi(Q_{\theta,\gamma,\alpha,\beta}\mathbb{Z}^{2})\varphi_{\frac{1}{\gamma}}\rangle\right\|_{\ell^{2}}$ $\asymp\left\|\langle f_{\gamma},\pi(R_{\theta}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$ by (7)
$=\left\|\langle\mathcal{F}^{(\theta)}f_{\gamma},\mathcal{F}^{(\theta)}\pi(R_{\theta}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$ by Proposition 1 (i)
$=\left\|\langle\mathcal{F}^{(\theta)}f_{\gamma},\pi(D_{\alpha,\beta}\mathbb{Z}^{2})\mathcal{F}^{(\theta)}\varphi\rangle\right\|_{\ell^{2}}$ by (8)
$=\left\|\langle\mathcal{F}^{(\theta)}f_{\gamma},\pi(D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$ by Proposition 1 (ii)
$\asymp\left\|\mathcal{F}^{(\theta)}f_{\gamma}\right\|_{L^{2}(\mathbb{R})}$ by maximality of $F(\varphi)$
$=\left\|f_{\gamma}\right\|_{L^{2}(\mathbb{R})}$ by Proposition 1 (i)
$=\left\|f\right\|_{L^{2}(\mathbb{R})},$ by (3)
which shows that $\mathcal{G}(\varphi_{1/\gamma},Q_{\gamma,\theta,\alpha,\beta}\mathbb{Z}^{2})$ is indeed a Gabor frame. ∎
Theorem 1.
If
$\lambda^{\prime}=-(\lambda+1/\lambda),\quad\gamma=\sqrt{\lambda^{2}+1},$ (9)
then the frame set of $h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{1/\gamma})$ is maximal.
Proof.
Let $(\alpha,\beta)$ be an arbitrary pair satisfying $\alpha,\beta>0$ and $\alpha\beta<1$. Set $D_{\alpha,\beta}=\operatorname{diag}(\beta,\alpha)$.
Let $U_{\lambda},L_{\lambda^{\prime}}$ be as defined in Lemma 1. Then for any $f\in L^{2}(\mathbb{R})$, straightforward computation by Lemma 1 (ii) and (iii) shows that
$\left\|\langle f,\pi(D_{\alpha,\beta}\mathbb{Z}^{2})\left(h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{1/\gamma})\right)\rangle\right\|_{\ell^{2}}$ $=\left\|\langle f,h_{\lambda}\cdot\pi(U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})(h_{\lambda^{\prime}}*\varphi_{1/\gamma})\rangle\right\|_{\ell^{2}}$
$=\left\|\langle h_{-\lambda}\cdot f,\pi(U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})(h_{\lambda^{\prime}}*\varphi_{1/\gamma})\rangle\right\|_{\ell^{2}}$
$=\left\|\langle h_{-\lambda}\cdot f,h_{\lambda^{\prime}}*\left(\pi(L_{\lambda^{\prime}}U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi_{1/\gamma}\right)\rangle\right\|_{\ell^{2}}$
$=\left\|\langle h_{-\lambda^{\prime}}*(h_{-\lambda}\cdot f),\pi(L_{\lambda^{\prime}}U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi_{1/\gamma}\rangle\right\|_{\ell^{2}}.$
By Lemma 1 (i) we have
$\left\|h_{-\lambda^{\prime}}*(h_{-\lambda}\cdot f)\right\|_{L^{2}(\mathbb{R})}=|\lambda^{\prime}|^{-\frac{1}{2}}\left\|f\right\|_{L^{2}(\mathbb{R})}.$
Therefore for the frame set of $h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{1/\gamma})$ to be maximal, we need to show that $\mathcal{G}(\varphi_{1/\gamma},L_{\lambda^{\prime}}U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})$ is a frame for all pairs of $(\alpha,\beta)$ with $\alpha,\beta>0$ and $\alpha\beta<1$.
Note that
$L_{\lambda^{\prime}}U_{\lambda}=\begin{pmatrix}1&0\\ \frac{1}{\lambda^{\prime}}&1\end{pmatrix}\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix}=\begin{pmatrix}1&\lambda\\ \frac{1}{\lambda^{\prime}}&\frac{\lambda}{\lambda^{\prime}}+1\end{pmatrix}$. (10)
If (9) holds, then the inner product between the two rows in $L_{\lambda^{\prime}}U_{\lambda}$ is
$1\cdot\frac{1}{\lambda^{\prime}}+\lambda\cdot(\frac{\lambda}{\lambda^{\prime}}+1)=\frac{1}{\lambda^{\prime}}(1+\lambda^{2}+\lambda\lambda^{\prime})=\frac{1}{\lambda^{\prime}}\left(1+\lambda^{2}-\lambda(\lambda+\frac{1}{\lambda})\right)=0,$
i.e., rows in $L_{\lambda^{\prime}}U_{\lambda}$ become orthogonal. Observe also that
$\det(L_{\lambda^{\prime}}U_{\lambda})=\det(L_{\lambda^{\prime}})\det(U_{\lambda})=1.$
Now since any $2\times 2$ real unitary matrix with determinant positive one is a rotation matrix of the form in (5), there must exist some some $\theta\in 2\pi\mathbb{T}$ so that we have
$L_{\lambda^{\prime}}U_{\lambda}=D_{\gamma}R_{\theta},$ (11)
for $D_{\gamma}=\operatorname{diag}(1/\gamma,\gamma)$ where
$\frac{1}{\gamma}=\frac{1}{\sqrt{1+\lambda^{2}}},\quad\gamma=\frac{1}{\sqrt{(\frac{1}{\lambda^{\prime}})^{2}+(\frac{\lambda}{\lambda^{\prime}}+1)^{2}}}=\sqrt{\lambda^{2}+1}$
are reciprocals of norms of the two rows in $L_{\lambda^{\prime}}U_{\lambda}$ respectively, which also comply with the value of $\gamma$ in (9). The two entries in the main diagonal of $D_{\gamma}$ must be reciprocals of each other since
$\det D_{\gamma}=\det(L_{\lambda^{\prime}}U_{\lambda}R_{-\theta})=\det(L_{\lambda^{\prime}})\det(U_{\lambda})\det(R_{-\theta})=1.$
Finally we set
$Q_{\gamma,\theta,\alpha,\beta}=D_{\gamma}R_{\theta}D_{\alpha,\beta}.$
Then we have
$\mathcal{G}(\varphi_{\frac{1}{\gamma}},L_{\lambda^{\prime}}U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})=\mathcal{G}(\varphi_{\frac{1}{\gamma}},Q_{\gamma,\theta,\alpha,\beta}\mathbb{Z}^{2}).$
By Lemma 2 we see that it is indeed a frame. ∎
3.3 Maximality: the general case
Lemma 3.
For any $\lambda,\lambda^{\prime}\in\mathbb{R}\setminus\{0\}$ and any $\gamma>0$, there is a complex number $s\in\mathbb{C}$ depending on $\lambda,\lambda^{\prime},\gamma$ such that
$h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{\frac{1}{\gamma}})=s\cdot h_{v+\lambda}\cdot\varphi_{\sqrt{u}},$
where
$u=\frac{\gamma^{2}}{\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2}},\quad v=\frac{\frac{1}{\lambda^{\prime}}}{\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2}}.$
Proof.
Recall the formula (see [16, A-7]) that if $\alpha,\beta\in\mathbb{R}$, $\alpha+\beta i\neq 0$ and $\alpha\geq|\beta|$, then
$\mathcal{F}(e^{-\pi(\alpha+\beta i)^{2}x^{2}})(\xi)=\frac{1}{\alpha+\beta i}e^{-\frac{\pi\xi^{2}}{(\alpha+\beta i)^{2}}}.$ (12)
Observe that
$h_{-\frac{1}{\lambda^{\prime}}}(\xi)\cdot\varphi_{\gamma}(\xi)=e^{\frac{\pi i\xi^{2}}{\lambda^{\prime}}}\cdot e^{-\pi\gamma^{2}\xi^{2}}=e^{-\pi\xi^{2}(\gamma^{2}-\frac{1}{\lambda^{\prime}}i)}=e^{\frac{-\pi\xi^{2}\left(\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2}\right)}{\gamma^{2}+\frac{1}{\lambda^{\prime}}i}}=e^{-\frac{\pi\xi^{2}}{u+vi}}.$
Clearly $u>0$. Thus $u+vi$ is in the right half complex plane. Hence there exists some $\eta=\alpha+\beta i\neq 0$ with $\alpha\geq|\beta|$ (i.e., the complex angle of $\eta$ is between $\pi/4$ and $\pi/4$) so that $\eta^{2}=u+vi$. It follows from (6) and (12) that
$(h_{\lambda^{\prime}}*\varphi_{\frac{1}{\gamma}})(x)=\mathcal{F}^{-1}(\hat{h}_{\lambda^{\prime}}\cdot\hat{\varphi}_{\frac{1}{\gamma}})(x)=\mathcal{F}^{-1}(c_{\lambda^{\prime}}h_{-\frac{1}{\lambda^{\prime}}}\cdot\gamma\varphi_{\gamma})(x)=c_{\lambda^{\prime}}\gamma\eta\cdot e^{-\pi(u+vi)x^{2}},$
where $c_{\lambda^{\prime}}$ is as defined in (6). Setting $s=c_{\lambda^{\prime}}\gamma\eta$, we obtain
$\left(h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{1/\gamma})\right)(x)=s\cdot e^{-\pi\left(u+(v+\lambda)i\right)x^{2}}=s\cdot h_{v+\lambda}(x)\cdot\varphi_{\sqrt{u}}(x).$
Plugging in the condition (9) of Theorem 1 into the imaginary part of the exponent above we get
$v+\lambda=\frac{\frac{1}{\lambda^{\prime}}}{\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2}}+\lambda$ $=\frac{-\frac{1}{\lambda+\frac{1}{\lambda}}}{(1+\lambda^{2})^{2}+(\frac{1}{\lambda+\frac{1}{\lambda}})^{2}}+\lambda$
$=\frac{-\frac{\lambda}{\lambda^{2}+1}}{(\lambda^{2}+1)^{2}+(\frac{\lambda}{\lambda^{2}+1})^{2}}+\lambda$
$=\frac{-\lambda(\lambda^{2}+1)}{(\lambda^{2}+1)^{4}+\lambda^{2}}+\lambda$
$=\lambda\cdot\frac{(\lambda^{2}+1)^{4}-1}{(\lambda^{2}+1)^{4}+\lambda^{2}}\neq 0,$
where the final inequality holds since $\lambda\neq 0$. This indicates that under the condition (9) of Theorem 1, $h_{\lambda}\cdot(h_{\lambda^{\prime}}*\varphi_{\frac{1}{\gamma}})=s\cdot h_{\lambda+v}\cdot\varphi_{\sqrt{u}}$ is actually a chirped Gaussian. Comparing with the characterization in (1) we see that it is indeed a new type of window with maximal frame sets.
So what if we make the columns in (10) orthogonal instead? This corresponds to the condition that $\lambda(\lambda^{\prime}+1/\lambda^{\prime})=-1$ and $\gamma=1$, and in such case we will have
$v+\lambda=\frac{\frac{1}{\lambda^{\prime}}}{1+(\frac{1}{\lambda^{\prime}})^{2}}-\frac{1}{\lambda^{\prime}+\frac{1}{\lambda^{\prime}}}=\frac{\lambda^{\prime}}{(\lambda^{\prime})^{2}+1}-\frac{\lambda^{\prime}}{(\lambda^{\prime})^{2}+1}=0,$
which means it is a trivial case since it produces nothing but merely a dilated Gaussian. The same happens if we replace the value of $\gamma$ in (9) with $1/\sqrt{\lambda^{2}+1}$.
Theorem 2.
The chirped Gaussian $h_{r}\cdot\varphi_{\sqrt{u}}$ has maximal frame set for any $r\in\mathbb{R}\setminus\{0\}$ and any $u>0$.
Proof.
For each $\lambda\in\mathbb{R}\setminus\{0\}$, there exists a pair $\lambda^{\prime},\gamma$ that satisfy (9) in Theorem 1 and in turn a pair $u,v$ satisfying Lemma 3 so that with $r=v+\lambda$ the chirped Gaussian $h_{r}\cdot\varphi_{\sqrt{u}}$ has maximal set.
For each $\gamma>0$, the dilated function $h_{\gamma^{2}r}\cdot\varphi_{\gamma\sqrt{u}}$ will also have maximal set since (4) indicates that $\mathcal{G}(h_{\gamma^{2}r}\cdot\varphi_{\gamma\sqrt{u}},\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is a frame if and only if $\mathcal{G}(h_{r}\cdot\varphi_{\sqrt{u}},\alpha\gamma\mathbb{Z}\times\beta/\gamma\mathbb{Z})$ is so, and the latter indeed is a frame by the maximality of $F(h_{r}\cdot\varphi_{\sqrt{u}})$.
The following ratio of parameters (of subscripts of $h$ and $\varphi$)
$\frac{\gamma^{2}r}{(\gamma\sqrt{u})^{2}}=\frac{r}{u},$
stays invariant upon dilation. Therefore to show that $h_{r}\cdot\varphi_{\sqrt{u}}$ has maximal frame set for any $r\in\mathbb{R}\setminus\{0\}$ and any $u>0$, it suffices for us to show that the function
$G(\lambda)=\frac{r}{u}=\frac{\lambda+v}{u},\quad\lambda\in\mathbb{R}\setminus\{0\},$
is surjective onto $\mathbb{R}\setminus\{0\}$. Direct computation shows that
$G(\lambda)=\frac{\lambda+v}{u}$ $=\frac{\lambda+\frac{\frac{1}{\lambda^{\prime}}}{\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2}}}{\frac{\gamma^{2}}{\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2}}}$
$=\frac{\lambda(\gamma^{4}+(\frac{1}{\lambda^{\prime}})^{2})+\frac{1}{\lambda^{\prime}}}{\gamma^{2}}$
$=\frac{1}{\gamma^{2}}\cdot\left(\lambda(\lambda^{2}+1)^{2}+\lambda(\frac{1}{\lambda+\frac{1}{\lambda}})^{2}-\frac{1}{\lambda+\frac{1}{\lambda}}\right)$
$=\frac{1}{\lambda^{2}+1}\cdot\left(\lambda(\lambda^{2}+1)^{2}+\lambda(\frac{\lambda}{\lambda^{2}+1})^{2}-\frac{\lambda}{\lambda^{2}+1}\right)$
$=\frac{\lambda}{\lambda^{2}+1}\cdot\left((\lambda^{2}+1)^{2}+\frac{\lambda^{2}}{(\lambda^{2}+1)^{2}}-\frac{1}{\lambda^{2}+1}\right)$
$=\frac{\lambda}{\lambda^{2}+1}\cdot\left(\lambda^{4}+2\lambda^{2}+\frac{\lambda^{2}}{(\lambda^{2}+1)^{2}}+\frac{\lambda^{2}}{\lambda^{2}+1}\right).$
Thus $G(\lambda)$ is a continuous function with
$\lim_{\lambda\to-\infty}G(\lambda)=-\infty,\quad\lim_{\lambda\to+\infty}G(\lambda)=+\infty.$
And $G(\lambda)=0$ only if $\lambda=0$. Therefore by the intermediate value property it is indeed surjective onto $\mathbb{R}\setminus\{0\}$. ∎
Since $|h_{r}\cdot\varphi_{\sqrt{u}}|=|\varphi_{\sqrt{u}}|$ it is also clear that $h_{r}\cdot\varphi_{\sqrt{u}}\in W_{0}(\mathbb{R})$. Then the Amalgam Balian-Low theorem excludes the critical case $\alpha\beta=1$ from its frame set.
Theorem 3.
Let $\gamma>0$ and $Q\in\mathbb{R}^{2\times 2}$. Then $\mathcal{G}(\varphi_{\gamma},Q\mathbb{Z}^{2})$ is a frame if and only if $0<|\det Q|<1$.
Proof.
We first show the sufficiency. Without loss of generality let us assume that $\det Q$ is positive. Otherwise consider Q-Q equivalently. We may also assume $\gamma=1$. Otherwise let
$Q^{\prime}=\operatorname{diag}(1/\gamma,\gamma)Q.$
Then (4) indicates that $\mathcal{G}(\varphi_{\gamma},Q\mathbb{Z}^{2})$ is a frame if and only if $\mathcal{G}(\varphi,Q^{\prime}\mathbb{Z}^{2})$ is a frame. And it is clear that $\det Q^{\prime}=\det Q$. Thus we can then just replace $\varphi_{\gamma}$ and $Q$ with $\varphi$ and $Q^{\prime}$, respectively.
Now suppose that $Q=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$ with $0<\det Q<1$. We apply a QR decomposition followed by a column scaling to get
$Q$ $=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
$=\begin{pmatrix}\frac{a}{\sqrt{a^{2}+c^{2}}}&-\frac{c}{\sqrt{a^{2}+c^{2}}}\\ \frac{c}{\sqrt{a^{2}+c^{2}}}&\frac{a}{\sqrt{a^{2}+c^{2}}}\end{pmatrix}\begin{pmatrix}\sqrt{a^{2}+c^{2}}&\frac{ab+cd}{\sqrt{a^{2}+c^{2}}}\\ 0&\frac{ad-bc}{\sqrt{a^{2}+c^{2}}}\end{pmatrix}$
$=\underbrace{\begin{pmatrix}\frac{a}{\sqrt{a^{2}+c^{2}}}&-\frac{c}{\sqrt{a^{2}+c^{2}}}\\ \frac{c}{\sqrt{a^{2}+c^{2}}}&\frac{a}{\sqrt{a^{2}+c^{2}}}\end{pmatrix}}_{=R_{\theta}}\underbrace{\begin{pmatrix}1&\frac{ab+cd}{ad-bc}\\ 0&1\end{pmatrix}}_{=U_{\lambda}}\underbrace{\begin{pmatrix}\sqrt{a^{2}+c^{2}}&0\\ 0&\frac{ad-bc}{\sqrt{a^{2}+c^{2}}}\end{pmatrix}}_{=D_{\alpha,\beta}}.$
The roles of notations $R_{\theta},U_{\lambda},D_{\alpha,\beta}$ are as given in Proposition 1, Lemma 1 and Lemma 2. In particular, we have
$\alpha\beta=\sqrt{a^{2}+c^{2}}\cdot\frac{ad-bc}{\sqrt{a^{2}+c^{2}}}=ad-bc=\det Q<1.$
Consequently for any $f\in L^{2}(\mathbb{R})$ we have
$\left\|\langle f,\pi(Q\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$ $=\left\|\langle f,\pi(R_{\theta}U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$
$=\left\|\langle\mathcal{F}^{(\theta)}f,\mathcal{F}^{(\theta)}\pi(R_{\theta}U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$ by Proposition 1(i)
$=\left\|\langle\mathcal{F}^{(\theta)}f,\pi(U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\mathcal{F}^{(\theta)}\varphi\rangle\right\|_{\ell^{2}}$ by Proposition 1(iii)
$=\left\|\langle\mathcal{F}^{(\theta)}f,\pi(U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$ by Proposition 1(ii)
$=\left\|\langle h_{\lambda}\cdot\mathcal{F}^{(\theta)}f,h_{\lambda}\cdot\pi(U_{\lambda}D_{\alpha,\beta}\mathbb{Z}^{2})\varphi\rangle\right\|_{\ell^{2}}$
$=\left\|\langle h_{\lambda}\cdot\mathcal{F}^{(\theta)}f,\pi(D_{\alpha,\beta}\mathbb{Z}^{2})\left(h_{\lambda}\cdot\varphi\right)\rangle\right\|_{\ell^{2}}$ by Lemma 2
$\asymp\left\|h_{\lambda}\cdot\mathcal{F}^{(\theta)}f\right\|_{L^{2}(\mathbb{R})}$ by Theorem 2
$=\left\|\mathcal{F}^{(\theta)}f\right\|_{L^{2}(\mathbb{R})}$
$=\left\|f\right\|_{L^{2}(\mathbb{R})},$ by Proposition 1(i)
which indicates that $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ is indeed a frame.
For the necessity, that $0<|\det Q|\leq 1$ is necessary for $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ to be a frame is asserted in [10, Theorem 10]. For the critical case of $1=|\det Q|=|\alpha\beta|$ we see that the above derivation actually indicates that $\mathcal{G}(\varphi,Q\mathbb{Z}^{2})$ is a frame if and only if $\mathcal{G}(h_{\lambda}\cdot\varphi,D_{\alpha,\beta}\mathbb{Z}^{2})$ is so. But $h_{\lambda}\cdot\varphi\in W_{0}(\mathbb{R})$. Thus by the Amalgam Balian-Low theorem it can not be a frame if $|\alpha\beta|=1$. ∎
4 Zeros of Zak transforms of chirped Gaussians
4.1 The Zak transform and the theta function
The Zak transform of $f\in L^{2}(\mathbb{R})$ is a function in $\mathbb{R}^{2}$ defined as
$Zf(t,\omega)=\sum_{k\in\mathbb{Z}}f(t-k)e^{2\pi ik\omega}.$
The unit square
$S=[0,1)\times[0,1)$
is called the fundamental domain of $Z$ since not only that $Z$ is an isometry from $L^{2}(\mathbb{R})$ to $L^{2}(S)$ (it maps the orthonormal Gabor basis element $M_{j}T_{k}\chi_{[0,1)}$ to the orthonormal Fourier basis element $e^{2\pi i(jt+k\omega)}$) but also it has the quasi-periodicity that
$Zf(t-j,\omega)=e^{-2\pi ij\omega}Z(t,\omega),\quad Zf(t,\omega-k)=Z(t,\omega),\quad\forall j,k\in\mathbb{Z},$
which indicates that the values of $Zf$ on $\mathbb{R}^{2}$ is completely determined by its values on $S$. See [6, Chapter 8] for more of its properties.
We shall call $(t_{0},\omega_{0})$ a simple zero of $Zf$ if $Zf(t_{0},\omega_{0})=0$ and there is a neighbourhood $\Omega$ of $(t_{0},\omega_{0})$ and a constant $C$ so that
$\frac{|Zf(t,\omega)|}{\sqrt{(t-t_{0})^{2}+(\omega-\omega_{0})^{2}}}\geq C,$
holds for all $(t,\omega)\in\Omega$.
So far all functions that found to have maximal frame sets possess only one simple zero in their fundamental domains, with the one sided exponential (which is not a continuous Wiener function) being the only exception: its Zak transform has no zero. We would like to give an exposition in the appendix (as they are not directly related to this paper) of known facts about zeros of the Zak transform since they seem to be playing some mysterious roles in the analysis of frame sets.
Theta functions are elliptic analogs of exponential functions. For convenience we will use the following form of the theta function, which is consist with the convention and results in [2]:
$\Theta(z,q)=\sum_{k\in\mathbb{Z}}q^{k^{2}}z^{k},\quad z\in\mathbb{C}\setminus\{0\},q\in\mathbb{C},|q|<1.$
This form is closest to the usual definition of
$\Theta_{3}(z,q)=\sum_{k\in\mathbb{Z}}q^{k^{2}}e^{2kiz},$
and can be easily translated into each other. As in [2], we also imposed the extra condition $|q|<1$. Otherwise the propositions below will hinge on absolute convergence of $\sum_{k}|q|^{k}$ and $\prod_{k}|1-q^{k}|$.
This form of the theta function is closely related to the Zak transform of Gaussians (see e.g.,[15]). In this article the following decomposition is relevant ([2, Lemma 2.1], which is also available in [27, Chapter 21.3, p.469]):
$\Theta(z,q)=\left(\prod_{j=1}^{\infty}(1-q^{2j})\right)\cdot\left(\prod_{k=0}^{\infty}(1+q^{2k+1}z)(1+q^{2k+1}z^{-1})\right).$
From this decomposition we immediately obtain that ([2, Corollary 2.2])
Proposition 2.
$\Theta(z,q)=0$ is only attained at $z=-q^{m}$ where $m$ can be any odd integer.
If we fix $q$, take $z=e^{2ix}$ for $x\in\mathbb{C}$ and set
$H(x)=\Theta(e^{2ix},q),$
then we have ([27, Chapter 21.3, p.469])
Proposition 3.
$H(x)$ is an entire function on $\mathbb{C}$ and all zeros of $H(x)$ are simple zeros.
4.2 Location, uniqueness, and simplicity of zeros
Theorem 4.
$Z(h_{\lambda}\cdot\varphi_{\gamma})$ has a unique simple zero at the center $(1/2,1/2)$ of the unit square $S$ for any $\lambda\in\mathbb{R}\setminus\{0\}$ and any $\gamma>0$.
Proof.
We readily compute
$Z(h_{\lambda}\cdot\varphi_{\gamma})(t,\omega)$ $=\sum_{k\in\mathbb{Z}}e^{-\pi i\lambda(t-k)^{2}}\cdot e^{-\pi\gamma^{2}(t-k)^{2}}\cdot e^{2\pi ik\omega}$
$=\sum_{k\in\mathbb{Z}}e^{-\pi i\lambda t^{2}}\cdot e^{2\pi i\lambda tk}\cdot e^{-\pi i\lambda k^{2}}\cdot e^{-\pi\gamma^{2}t^{2}}\cdot e^{2\pi\gamma^{2}tk}\cdot e^{-\pi\gamma^{2}k^{2}}\cdot e^{2\pi ik\omega}$
$=e^{-\pi(\gamma^{2}+\lambda i)t^{2}}\sum_{k\in\mathbb{Z}}e^{-\pi(\gamma^{2}+\lambda i)k^{2}}\cdot e^{2\pi\left(\gamma^{2}t+(\omega+\lambda t)i\right)k}$
$=e^{-\pi(\gamma^{2}+\lambda i)t^{2}}\Theta(z,q),$
where
$z=e^{2\pi\left(\gamma^{2}t+(\omega+\lambda t)i\right)},\quad q=e^{-\pi(\gamma^{2}+\lambda i)}.$
This is well defined since $|q|<1$ and $z\neq 0$ is obvious. By Proposition 2, zeros are only attained at $z=-q^{m}$ for odd integers $m$. Taking absolute values at both sides we get
$e^{2\pi\gamma^{2}t}=|z|=|q^{m}|=e^{-\pi\gamma^{2}m}.$
For $t\in[0,1)$ the equation can only hold at $t=1/2$, in which case m=1m=-1. Plugging this back in we obtain
$e^{2\pi\left(\frac{\gamma^{2}}{2}+(\omega+\frac{\lambda}{2})i\right)}=z=-q^{-1}=-e^{\pi(\gamma^{2}+\lambda i)},$
i.e., $e^{2\pi\omega i}=-1$, and for $\omega\in[0,1)$ this is only attained at $\omega=1/2$. Therefore $(1/2,1/2)$ is the only zero of $Z(h_{\lambda}\cdot\varphi_{\gamma})$ in $S$.  [^1] Since $h_{\lambda}\cdot\varphi$ is an even function, the existence of a zero at the center of $S$ can also be asserted by Proposition 4 (c) in the appendix, but the uniqueness of this zero is non-trivial.
Finally since $q$ does not depend on $t,\omega$, by Proposition 3 we see that the zero is simple with respect to the complex variable
$x=\frac{2\pi\left(\gamma^{2}t+(\omega+\lambda t)i\right)}{2i}=\pi(\omega+\lambda t-\gamma^{2}ti).$
We set
$z=\begin{pmatrix}\omega\\ t\end{pmatrix},\quad z_{0}=\begin{pmatrix}1/2\\ 1/2\end{pmatrix},\quad U=\begin{pmatrix}1&\lambda\\ 0&\gamma^{2}\end{pmatrix},\quad z_{1}=U^{-1}z_{0},$
where $U^{-1}$ exists since $\gamma\neq 0$. Then there are constants $C,C^{\prime}$ so that we have ($\|\cdot\|$ is the Euclidean norm)
$|Z(h_{\lambda}\cdot\varphi_{\gamma})(t,\omega)|^{2}$ $\geq C|x-\frac{1}{2}-\frac{1}{2}i|^{2}$
$=C(\omega+\lambda t-\frac{1}{2})^{2}+(\gamma^{2}t-\frac{1}{2})^{2}$
$=C\|Uz-z_{0}\|^{2}$
$=C\|U(z-z_{1})\|^{2}$
$\geq C^{\prime}\|z-z_{1}\|^{2},$
[…]
$\|z-z_{1}\|^{2}\geq C^{\prime\prime}\|z-z_{0}\|^{2},\quad\forall z\in\Omega.$
Together we can conclude that $(1/2,1/2)$ is indeed a simple zero of $Z(h_{\lambda}\cdot\varphi_{\gamma})$. ∎
Appendix: Facts about zeros of the Zak transform
In this section for the maximality of $F(g)$, instead of asking whether $(\alpha,\beta)\in F(g)$ holds for all $\alpha,\beta>0$ with $\alpha\beta<1$, we consider equivalently whether $(\alpha,1)\in F(g_{\gamma})$ holds for all $0<\alpha<1$ and all $\gamma\in\mathbb{R}^{+}$. This is a just simple consequence of (4) since $\mathcal{G}(g,\alpha\mathbb{Z}\times\beta\mathbb{Z})$ is a frame if and only if $\mathcal{G}(g_{1/\beta},\alpha\beta\mathbb{Z}\times\mathbb{Z})$ is.
The Zak transform method is particularly effective in analyzing frame sets of the critical case and other integer oversampling cases since in such cases it diagonalizes the frame operator and makes it a multiplication operator [6, Corollary 8.3.2, Equation 8.21], i.e., if $\alpha=1/n$ with $n\in\mathbb{N}$, then $\mathcal{G}(g_{\gamma},\alpha\mathbb{Z}\times\mathbb{Z})$ is a frame if and only if
$\sum_{j=0}^{n-1}|Zg_{\gamma}(t-\frac{j}{n},\omega)|^{2}\asymp 1$
holds for almost every $t,\omega$. Therefore in these cases locations of zeros of $Zg_{\gamma}$ play a prominent role, since as long as $Zg_{\gamma}$ is continuous, then its essential lower boundedness on $S$ (which by quasi-periodicity is equivalent to its global essential lower boundedness) is equivalent to having no zero in $S$.
The other cases are more complicated. But in general they can be characterized as ([11], also stated as [8, Theorem 2.3]): $\mathcal{G}(g_{\gamma},\alpha\mathbb{Z}\times\mathbb{Z})$ is a frame if and only if
$\sum_{k\in\mathbb{Z}}\left|\int_{0}^{1}f(\omega)Zg_{\gamma}(t-k\alpha,\omega)d\omega\right|^{2}\asymp\|f\|_{L^{2}([0,1))}^{2}$ (13)
holds for all $f\in L^{2}([0,1))$ and all $t\in[0,1)$.
To understand how the estimates change upon dilation is often a major difficulty since it is hard to relate and control $Zg_{\gamma}$ through $Zg$. Another issue is that unlike in the critical case or in other integer oversampling cases, here merely lifting the lower bound of $|Zg_{\gamma}|$ does not seem to help. Indeed, for an individual $g_{\gamma}$, if we fix $t$, then for each $k\in\mathbb{Z}$, the function $h_{k}^{(t)}(\omega)=Zg_{\gamma}(t-k\alpha,\omega)$ is a univariate function of $\omega$, and the sum in (13) becomes $\sum_{k}|\langle f,\overline{h_{k}^{(t)}}\rangle_{L^{2}([0,1])}|^{2}$. Therefore (13) is actually equivalent to requiring $\{h_{k}^{(t)}\}_{k}$ to be a frame of $L^{2}([0,1))$. Hence even if each $|h_{k}^{(t)}|$ is away from 0, together they can still be collided together for some $f$ to be orthogonal to all of them to make the sum in (13) vanish.
Despite that, as mentioned, mysteriously so far all functions that found to have maximal frame sets possess at most one simple zero in $S$. Hence we would like to provide an overview of what is known (other than Theorem 4) about existence, simplicity and uniqueness of zeros of Zak transforms.
Proposition 4.
Existence:
  1. (a)
    If $Zf$ is continuous, then it has a zero in $S$.
  2. (b)
    If $f$ is real valued, then zeros of $Zf$ in $S$ are symmetric with respect to the line $\omega=1/2$. If in addition $Zf$ is continuous, then $Zf$ must have a zero on the line $\omega=1/2$.
  3. (c)
    If $f$ is an even function, then $Zf(1/2,1/2)=0$.
  4. (d)
    If $f$ is an odd function, then $Zf(0,0)=Zf(1/2,0)=Zf(0,1/2)=0$.
  5. (e)
    If $f,\hat{f}\in W(\mathbb{R})$ and $f$ is an eigenfunction of $\mathcal{F}$ whose eigenvalue is not i-i, then $Zf(1/2,1/2)=0$.
  6. (f)
    If $f\in L^{1}(\mathbb{R})$ is non-negative, supported on and strictly decreasing on $0,+\infty)$, then there is some constant $A$ so that $|Zf|>A>0$ holds everywhere.
Uniqueness and simplicity:
  1. (1)
    If $f$ is even, continuous, and can be written as $g+T_{1}g$ on $0,+\infty)$ for some strictly convex, non-negative $g\in L^{1}(\mathbb{R})$, then $Zf$ has a unique zero in $S$.
  2. (2)
    If $f\in L^{1}(\mathbb{R})$ is even, continuous, non-negative and strictly convex on $0,+\infty)$, then $Zf$ has no zero on $S\setminus\{(t,\omega):t=1/2\}$.
  3. (3)
    If $f$ is an exponential $B$-spline of finite order or a totally positive function (excluding the one sided exponential and its translations) or the convolution of these two types of functions, then $Zf$ has a unique zero in $S$, the zero is simple and is on the line $\omega=1/2$.
(a) is a classical result, and we refer to [6, Lemma 8.4.2] for a proof and references therein for its origin. (f) is [14, Section 3.2.1].
(b) to (e) are mentioned at many places. But they are often mixed with stronger assumptions. We give a simple sketch here: for (b), if $f$ is real valued then
$\overline{Zf(t,\omega)}=\sum_{k\in\mathbb{z}}\overline{f(t-k)e^{2\pi ik\omega}}=\sum_{k\in\mathbb{z}}f(t-k)e^{-2\pi ik\omega}=Zf(t,-\omega)=Zf(t,1-\omega),$
where the last equality follows from the quasi-periodicity. This indicates that all zeros are symmetric with respect to the line $\omega=1/2$, and on this line we have that
$Zf(t,\frac{1}{2})=\sum_{k\in\mathbb{Z}}f(t-k)\cdot(-1)^{k},$
is also a real valued function of $t$, then $Zf$ being continuous implies it is also continuous. By the quasi-periodicity we have Zf(0,1/2)=Zf(1,1/2)Zf(0,1/2)=-Zf(1,1/2), thus leading to the existence of a zero for some $t\in[0,1)$ by the intermediate value property.
For (c), if $f$ is even, then
$Zf(t,\omega)=\sum_{k\in\mathbb{Z}}f(t-k)e^{2\pi ik\omega}=\sum_{k\in\mathbb{Z}}f(k-t)e^{2\pi ik\omega}.$
Replacing $k$ with j-j we obtain
$\sum_{k\in\mathbb{Z}}f(k-t)e^{2\pi ik\omega}=\sum_{j\in\mathbb{Z}}f(-j-t)e^{-2\pi ij\omega}=\sum_{j\in\mathbb{Z}}f(j+t)e^{-2\pi ij\omega}.$
Plugging in $(t,\omega)=(1/2,1/2)$ we see that both equations are signed sums of $f$ at half integers. But they have opposite signs, i.e., Zf(1/2,1/2)=Zf(1/2,1/2)Zf(1/2,1/2)=-Zf(1/2,1/2), which means it is 0.
For (d), if $f$ is odd, then similar as above we have
$Zf(t,\omega)=-\sum_{k\in\mathbb{Z}}f(k-t)e^{2\pi ik\omega}=\sum_{j\in\mathbb{Z}}f(j+t)e^{-2\pi ij\omega}.$
Plugging in $(t,\omega)=(1/2,0)$ we get sums of $f$ at half integers with opposite signs. Thus $Zf(1/2,0)=0$, and it is also clear from oddity by direct computation that $Zf(0,0)=Zf(0,1/2)=f(0)=0$.
For (e), suppose that $\lambda\neq-i$ is the eigenvalue. Then we simply apply the formula ([6, Proposition 8.2.2]) that
$Zf(t,\omega)=e^{2\pi it\omega}Z\hat{f}(\omega,-t),\quad f,\hat{f}\in W(\mathbb{R})$
followed by the quasi-periodicity to get
$Zf(t,\omega)=e^{2\pi it\omega}Z\hat{f}(\omega,-t)=\lambda e^{2\pi it\omega}Zf(\omega,-t)=\lambda e^{2\pi it\omega}Zf(\omega,1-t).$
Plugging in $t=\omega=1/2$ we obtain $Zf(1/2,1/2)=i\lambda Zf(1/2,1/2)$, and $i\lambda\neq 1$ since $\lambda\neq-i$, which shows (e).
(1) and (2) are in [14, Section 3.2.2-3.2.3]. (3) is established in [2] for Gaussians, in [17] for exponential $B$-splines of finite order and for finite type totally positive functions without the Gaussian factor, in [18] for infinite type totally positive functions without the Gaussian factor, and finally in [26] for totally positive functions with the Gaussian factor and for their convolutions with exponential $B$-splines.